What is the electric flux through one side of a cube that has a single point charge of -2 µC placed at its center? Hint: You do not need to integrate any equations to get the answer.

N·m2/C
Answer #1

As we know that

Since there are six identical faces to the cube, each face gets flux
Φ = Q / 6*εo = -2e-6C / 6*8.852e-12C²/N·m² = -3.766478343*10^4 V·m
Note that this is independent of the size of the cube.